# Climate Science and Control Engineering: Insights, Parallels, and Connections

### v2

@@ V-A Dynamic Climate Models 101: The Unforced EBM
$\leavevmode\nobreak\ C\frac{dT(t)}{dt}=S(1-\alpha)-\epsilon\sigma T(t)^{4}.$ (1)
@@ V-B Modeling Control Inputs and GHGs Emissions: Formulating a Forced EBM
$\leavevmode\nobreak\ C\frac{dT(t)}{dt}=S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))-\epsilon\sigma T(t)^{4}.$ (2)
@@
$\leavevmode\nobreak\ C\frac{dT(t)}{dt}=S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))-(\epsilon+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)})\sigma T(t)^{4},$ (3)
@@
If no explicit control is applied (i.e., $u(t)=0$) and emissions rise unchecked (i.e., $w(t)<0$), the system experiences increased warming due to enhanced radiative trapping. Conversely, effective control ($u(t)>0$) and stable or reduced emissions ($w(t)\approx 0$) can lead to stabilization or even gradual cooling.
@@
The parameters used in this simulation are: effective heat capacity $C=8\times 10^{8}\,\text{J/K/m}^{2}$, solar constant $S=1367.6\,\text{W/m}^{2}$, base albedo $\alpha=0.3$, and nominal emissivity $\epsilon=0.61$. The control input corresponds to an increase in albedo of $u(t)=0.2$, while the disturbance is modeled as a decrease in emissivity of $w(t)=-0.15$. The initial temperature is set to $T_{0}=288\,\text{K}$ (equivalent to $15^{\circ}$), and the simulation time span covers 30 years. Gaussian noise with a standard deviation of 10% is applied to the model dynamics, while the parameters $\alpha$ and $\epsilon$ are randomly perturbed with a standard deviation of approximately 1% around their nominal values to represent structural uncertainty. To maintain consistency with the rest of the tutorial, where temperature is presented in degrees Celsius, all plotted results are shifted from Kelvin to Celsius by subtracting 273.15. This makes the interpretation of the temperature trajectories more intuitive and aligned with conventional climate science practices.
@@
Having looked at the EBM from this perspective, some control-type questions can be posed such as: What input is needed to stabilize temperature at a target? How sensitive is the system to delays or uncertainty in the control action? These kinds of questions bring forward important concepts that connect climate modeling with insights from control engineering as discussed in Section VII.
@@ VI More Complex Dynamic Climate Model
$\displaystyle\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}+f\mathbf{k}\times\mathbf{v}$ $\displaystyle=-\frac{1}{\rho}\nabla p+\mathbf{F}_{\text{viscous}},\leavevmode\nobreak$ (4a)
$\displaystyle\frac{\partial T}{\partial t}+\mathbf{v}\cdot\nabla T$ $\displaystyle=Q_{T},\leavevmode\nobreak$ (4b)
$\displaystyle\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{v})$ $\displaystyle=0.\leavevmode\nobreak$ (4c)
@@
[…]
@@ VI-B Mathematical Interpretation of the Atmospheric Primitive Model
[…]
@@
      $\leavevmode\nobreak\ \frac{\partial T}{\partial t}=-\left(u\frac{\partial T}{\partial x}+v\frac{\partial T}{\partial y}+w\frac{\partial T}{\partial z}\right)+Q_{T}.$ (6)
@@
[…]
@@ VI-C Solving the Atmospheric Primitive Model
[…]
@@
[…]
@@
[…]
@@
Model Type Horizontal Resolution Vertical Layers Time Step Notes
@@
$\mathbf{v}^{n}:=\mathbf{v}(t_{n}),\quad\mathbf{T}^{n}:=\mathbf{T}(t_{n}),\quad\bm{\rho}^{n}:=\bm{\rho}(t_{n}),$
@@ VI-D State-Space Outlook on the Atmospheric Model
[…]
@@
[…]
Raising Control Engineering Questions. The models presented in (4) and (10) open up classical and modern control questions that are related to core control concepts, such as the following, pertaining to specifically observability and controllability.
@@
  • Observability and Sensor Placement: Which parts of the atmospheric state can be reconstructed from the available measurements? For specific kind of sensors, what are the unobservable subspaces in one specific climate model? What and how many sensors are required to ensure desirable observability metrics for any of the derived climate models? These questions are often posed in dynamic system sciences but are hardly navigated in climate research, in this particular system-theoretic context.
@@
  • Controllability, Reachability Analysis, and Control Node Placement: If hypothetical actuators are introduced to imitate controlled practices (e.g., geoengineering techniques) and policy-driven mitigation schemes, then mapped to present control vectors in (4) and (10), then how effective are these control measures at steering the system toward a more desirable climate trajectory? What are the resulting controllable subspaces and how can reachability analysis improve our understanding of the decade-long evolution of climate states? Instead of just simply running extended period simulations of climate models, can reachability analysis inform or predict reachable spaces given improved mitigation strategies and specific control actions? What is the impact of placing control nodes as actuators in specific geographic locations to thereby improve the trajectory of climate states? Can this be posed as an actuator placement problem?

### v3

@@ V-A Dynamic Climate Models 101: The Unforced EBM
$\penalty 10000\ C\frac{dT(t)}{dt}=S(1-\alpha(t))-\epsilon\sigma T(t)^{4}.$ (1)
This EBM captures how the global mean temperature evolves over time in response to the imbalance between absorbed solar energy and emitted infrared radiation. In doing so, it implicitly represents the combined thermal response of the Earth’s surface and atmosphere, abstracting away more detailed processes in the ocean, cryosphere, and biosphere.
@@ V-B Modeling Control Inputs and GHGs Emissions: Formulating a Forced EBM
$\penalty 10000\ C\frac{dT(t)}{dt}=S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))-\epsilon\sigma T(t)^{4}.$ (2)
@@
$\penalty 10000\ C\frac{dT(t)}{dt}=S(1-(\alpha+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}))-(\epsilon+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)})\sigma T(t)^{4},$ (3)
@@
If no explicit control is applied (i.e., ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}=0$) and emissions rise unchecked (i.e., ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)}<0$), the system experiences increased warming due to enhanced radiative trapping. Conversely, effective control (${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}>0$) and stable or reduced emissions (${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)}\approx 0$) can lead to stabilization or even gradual cooling.
@@
The parameters used in this simulation are: effective heat capacity $C=8\times 10^{8}\,\text{J/K/m}^{2}$, solar constant $S=1367.6\,\text{W/m}^{2}$, base albedo $\alpha=0.3$, and nominal emissivity $\epsilon=0.61$. The control input corresponds to an increase in albedo of ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u(t)}=0.2$, while the disturbance is modeled as a decrease in emissivity of ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)}=-0.15$. The initial temperature is set to $T_{0}=288\,\text{K}$ (equivalent to $15^{\circ}$C), and the simulation time span covers 30 years. Gaussian noise with a standard deviation of 10% is applied to the model dynamics, while the parameters $\alpha$ and $\epsilon$ are randomly perturbed with a standard deviation of approximately 1% around their nominal values to represent structural uncertainty. To maintain consistency with the rest of the tutorial, where temperature is presented in degrees Celsius, all plotted results are shifted from Kelvin to Celsius by subtracting 273.15. This makes the interpretation of the temperature trajectories more intuitive and aligned with conventional climate science practices.
@@
A more advanced EBM could also be derived considering the Earth as a grid with $m$ latitude and $n$ longitude points. Each cell in this grid of dimension $mn$ has its own dynamical model depicting surface temperature evolution $T_{ij}(t)$ in cell $(i,j)$ that basically extends (1)–(3) into the following:
$\displaystyle\small C_{ij}\frac{d{T}_{ij}(t)}{dt}=S_{ij}(t)\,\big(1-\alpha_{ij}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u_{ij}(t)}\big)\penalty 10000\$(4)
$\hskip 85.35826pt-\varepsilon_{ij}\,\sigma\,T_{ij}^{4}(t)+\mathcal{L}_{ij}(\{T_{kl}(t)\})+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F_{ij}(t)},$
where $i\in\{1,2,\ldots,m\}$ and $j\in\{1,2,\ldots,n\}$ define the latitude and longitude indices. The other quantities in (4) are defined similarly to (1)–(3). The major difference between the average EBM and the 2D model (4) is the function $\mathcal{L}$ which represents the diffusive heat exchange with neighboring cells. The diffusive term $\mathcal{L}_{ij}(\{T_{kl}\})$ accounts for horizontal heat transport between neighboring grid cells $(k,l)\in\mathcal{N}(i,j)$, with diffusion coefficients $\kappa_{ij,kl}$ which is typically approximated with a linear model, where the values for $\kappa$’s vary according to the considered assumptions and to the specific application area of the climate model. The quantity $F_{ij}(t)$ represents external forcing or unresolved subgrid effects, akin to the ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w(t)}$ in (3). This 2D EBM presents itself as a networked system of temperature dynamics $T_{ij}(t)$ where the coupling appears in the $\mathcal{L}_{ij}(\cdot)$ function (albeit the coupling is very sparse as every local temperature $T_{ij}(t)$ only relates to its adjacent temperatures in the grid), in contrast to the simpler single-dimension EBM which averages individual $mn$ quantities $C_{ij},S_{ij},\alpha_{ij},\epsilon_{ij}$ into single quantities. Later in this paper we show how more advanced models can be derived into controls-oriented state-space description.
Having looked at the basic EBM and its variants, some control-type questions can be posed such as: What input is needed to stabilize temperature at a target? How sensitive is the system to delays or uncertainty in the control action? These kinds of questions bring forward important concepts that connect climate modeling with insights from control engineering as discussed in Section VIII.
@@ VI More Complex Dynamic Climate Model
$\displaystyle\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}+f\mathbf{k}\times\mathbf{v}$ $\displaystyle=-\frac{1}{\rho}\nabla p+\mathbf{F}_{\text{viscous}},\penalty 10000\$(5a)
$\displaystyle\frac{\partial T}{\partial t}+\mathbf{v}\cdot\nabla T$ $\displaystyle=Q_{T},\penalty 10000\$(5b)
$\displaystyle\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{v})$ $\displaystyle=0.\penalty 10000\$(5c)
@@
In (5), the symbol $\times$ denotes the vector cross product operator. The operator $\nabla$ represents the spatial gradient. When applied to a scalar field $\phi$, it yields the vector of partial derivatives: $\nabla\phi=\left[\frac{\partial\phi}{\partial x},\frac{\partial\phi}{\partial y},\frac{\partial\phi}{\partial z}\right]^{\top}$. When applied to a vector field, $\nabla$ returns the Jacobian matrix of partial derivatives. The divergence operator $\nabla\cdot$ maps a vector field to a scalar; for example, $\nabla\cdot\mathbf{a}=\frac{\partial a_{x}}{\partial x}+\frac{\partial a_{y}}{\partial y}+\frac{\partial a_{z}}{\partial z}$ returns the net scalar outflow of $\mathbf{a}$.
@@ VI-B Mathematical Interpretation of the Atmospheric Primitive Model
    • $f\mathbf{k}\times\mathbf{v}$ is the Coriolis force, which represents the apparent force due to Earth’s rotation. The Coriolis parameter $f$ depends on latitude, and $\mathbf{k}=[0,0,1]^{\top}$ is the unit vector in the vertical direction. The cross product expands as $f\mathbf{k}\times\mathbf{v}=[-fv,\ fu,\ 0]^{\top}.$
@@
      $\penalty 10000\ \frac{\partial T}{\partial t}=-\left(u\frac{\partial T}{\partial x}+v\frac{\partial T}{\partial y}+w\frac{\partial T}{\partial z}\right)+Q_{T}.$ (7)
@@
      $\penalty 10000\ \begin{aligned} \frac{\partial\rho}{\partial t}=&-\Bigg(u\frac{\partial\rho}{\partial x}+v\frac{\partial\rho}{\partial y}+w\frac{\partial\rho}{\partial z}\\ &+\rho\left(\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}+\frac{\partial w}{\partial z}\right)\Bigg).\end{aligned}$ (8)
@@ VI-C Solving the Atmospheric Primitive Model
$\nabla\cdot$ [1/m] $\mathbb{R}^{N\times 3N}$ Divergence operator Maps vector fields $\rightarrow$ scalar fields. Used in mass conservation (continuity) equation
@@
Spatial Discretization (Semi-Discrete Form). The continuous spatial fields are replaced with grid-defined fields over $N$ nodes. Derivative operators such as $\nabla$ and $\nabla\cdot$ are approximated numerically using one of the following methods:
@@
$\displaystyle\begin{aligned} \frac{d}{dt}\begin{bmatrix}\mathbf{v}_{x}\\ \mathbf{v}_{y}\\ \mathbf{v}_{z}\end{bmatrix}&=-\begin{bmatrix}\mathbf{v}_{x}\odot\frac{\partial\mathbf{v}_{x}}{\partial x}+\mathbf{v}_{y}\odot\frac{\partial\mathbf{v}_{x}}{\partial y}+\mathbf{v}_{z}\odot\frac{\partial\mathbf{v}_{x}}{\partial z}\\ \mathbf{v}_{x}\odot\frac{\partial\mathbf{v}_{y}}{\partial x}+\mathbf{v}_{y}\odot\frac{\partial\mathbf{v}_{y}}{\partial y}+\mathbf{v}_{z}\odot\frac{\partial\mathbf{v}_{y}}{\partial z}\\ \mathbf{v}_{x}\odot\frac{\partial\mathbf{v}_{z}}{\partial x}+\mathbf{v}_{y}\odot\frac{\partial\mathbf{v}_{z}}{\partial y}+\mathbf{v}_{z}\odot\frac{\partial\mathbf{v}_{z}}{\partial z}\end{bmatrix}\\ &\quad-f\begin{bmatrix}-\mathbf{v}_{y}\\ \mathbf{v}_{x}\\ \mathbf{0}\end{bmatrix}-\bm{\rho}^{-1}\odot\begin{bmatrix}\frac{\partial\mathbf{p}}{\partial x}\\ \frac{\partial\mathbf{p}}{\partial y}\\ \frac{\partial\mathbf{p}}{\partial z}\end{bmatrix}+\begin{bmatrix}\mathbf{F}_{x}\\ \mathbf{F}_{y}\\ \mathbf{F}_{z}\end{bmatrix}\end{aligned}$ (10a)
$\displaystyle\begin{aligned} \frac{d\mathbf{T}}{dt}&=-\left(\mathbf{v}_{x}\odot\frac{\partial\mathbf{T}}{\partial x}+\mathbf{v}_{y}\odot\frac{\partial\mathbf{T}}{\partial y}+\mathbf{v}_{z}\odot\frac{\partial\mathbf{T}}{\partial z}\right)+\mathbf{Q}_{T}\end{aligned}$ (10b)
$\displaystyle\begin{aligned} \frac{d\bm{\rho}}{dt}&=-\left(\mathbf{v}_{x}\odot\frac{\partial\bm{\rho}}{\partial x}+\mathbf{v}_{y}\odot\frac{\partial\bm{\rho}}{\partial y}+\mathbf{v}_{z}\odot\frac{\partial\bm{\rho}}{\partial z}\right)\\ &\quad-\bm{\rho}\odot\left(\frac{\partial\mathbf{v}_{x}}{\partial x}+\frac{\partial\mathbf{v}_{y}}{\partial y}+\frac{\partial\mathbf{v}_{z}}{\partial z}\right)\end{aligned}$ (10c)
@@
$\mathbf{v}^{n}=\mathbf{v}(t_{n}),\quad\mathbf{T}^{n}=\mathbf{T}(t_{n}),\quad\bm{\rho}^{n}=\bm{\rho}(t_{n}),$
@@
Model Type Horizontal Resolution Vertical Layers Time Step Notes
@@ VI-D State-Space Outlook on the Atmospheric Model
The system is originally described by nonlinear PDEs governing momentum, thermodynamics, and mass conservation. Upon discretization in space and time (i.e., Eq. (11)), the system becomes a high-dimensional, nonlinear discrete-time dynamical system. The dimensionality of the resulting system is directly tied to the spatial resolution of the grid and the number of physical variables tracked at each grid point.
@@
The state vector includes grid-point values of wind velocity $\mathbf{v}$, temperature $\mathbf{T}$, and density $\bm{\rho}$ across the 3D model domain. Measurable outputs typically include surface or upper-air observations of temperature, wind speed, and pressure, obtained from satellites, radiosondes, or ground stations. In addition to states, some external forcing terms or boundary inputs, such as radiative heating, surface fluxes, or pressure gradients, may be estimated or prescribed from measurements or reanalysis data. Since the full state and all inputs are not directly observable, estimation and inference play a central role. These challenges are well-studied in the control engineering field, and their relevance to atmospheric modeling is further discussed in Section VIII.
Similarly to the EBM model (2), control inputs can be conceptually introduced to the atmospheric system to represent hypothetical or indirect interventions. In physical terms, these inputs would correspond to modifications of the external forcing terms (e.g., heat sources $Q_{T}$), boundary conditions, or parameterized sub-grid processes. Unlike in the EBM, where the input can directly modify surface albedo or emissivity in a scalar ODE, here the influence would be distributed in space and time across the 3D model grid. For example, injecting stratospheric aerosols would alter the radiative balance in the energy equation by increasing reflectivity and thus reducing incoming solar energy—a localized change in the heating term. Similarly, implementing large-scale land-use changes or reflective surfaces could alter surface properties and boundary conditions. Though these inputs are not part of standard weather or climate prediction workflows, they are useful constructs in a control-theoretic interpretation of the system: they represent degrees of influence over the model’s evolution, enabling the posing of questions such as: what form, location, or magnitude of intervention is needed to influence atmospheric temperature distributions or circulation patterns? These inputs can be thought of as distributed, constrained actuators that interface with the physical system, similar in spirit to how geoengineering or emission policies have been framed as control inputs in the EBM case.
@@
Raising Control Engineering Questions. The models presented in (5) and (11) open up classical and modern control questions that are related to core control concepts, such as the following, pertaining to specifically observability and controllability. For observability-related analyses, the following questions can be asked: Which parts of the atmospheric state $\bm{x}(t)$ can be reconstructed from the available measurements $\bm{y}(t)$? What and how many sensors are required to ensure desirable observability metrics for any of the derived climate models? These questions are often posed in dynamic system sciences but are hardly navigated in climate research.
@@
As for controllability analysises, the following questions can be posed. If hypothetical actuators are introduced to imitate controlled practices (e.g., geoengineering techniques) and policy-driven mitigation schemes, then mapped to present control vectors in (5) and (11), then how effective are these control measures at steering the system toward a more desirable climate trajectory? What are the resulting controllable subspaces and how can reachability analysis improve our understanding of the decade-long evolution of climate states? What is the impact of placing control nodes as actuators in specific geographic locations to thereby improve the trajectory of climate states? Can this be posed as an actuator placement problem?